A buffer consists of 0.65 M of the weak monoprotic base B and 0.50 M of...

  1. A buffer consists of 0.65 M of the weak monoprotic base B and 0.50 M of it salt BHCl. What is the pH of the buffer after 0.006 moles of KOH are added to 0.20 L of this buffer solution? Kb for the monoprotic base B is 1.3× 10‒3.
Answer #1

Given buffer contain weak base B and its conjugate acid BH + . K b of weak base is 1.3 \times 10 -3.

We have relation, pK b = - log K b

\therefore  pK b = - log ( 1.3 \times 10 -3. ) = 2.89

Now, calculate concentration of B & BH + in moles.

We have, Molarity = No. of moles of solute / volume of solution in L

\therefore No. of moles of solute = Molarity \times volume of solution in L

\therefore No. of moles of B = 0.65 mol / L \times 0.20 L = 0.13 mol

\therefore No. of moles of BH + = 0.5 mol / L \times 0.20 L = 0.10 mol

Consider reaction of KOH with buffer solution. KOH reacts with BHCl and produces water and base B.

BHCl + KOH \rightarrow B + H2O + KCl

Let' use ICE table.

Concentration ( moles ) BHCl KOH B
I 0.10 0.006 0.13
C - 0.006 -0.006 +0.006
E 0.094 0.000 0.136

We can calculate pOH of buffer solution by using Henderson's equation as shown below.

pOH = pK b + log [ salt] / [ Base ]

\therefore pOH = 2.89 + log [ BHCl ] / [ B]

\therefore pOH = 2.89 + log ( 0.094 mol / 0.20 L )  / ( 0.136 mol / 0.20 L )

p OH = 2.89 - 0.1604

p OH = 2.73

We have relation, p H + p OH = 14

\therefore p H = 14 - p OH = 14 - 2.73 = 11.27

ANSWER : pH of buffer solution after addition of 0.006 mol KOH = 11.27